OpenStudy (anonymous):

FINAL REVIEW HELP : sqrt{6}(7sqrt{12}- 10sqrt{3})

OpenStudy (anonymous):

√ 12 = √ (4 * 3 ) = ...

OpenStudy (anonymous):

\[12\sqrt{2}\] ?

OpenStudy (anonymous):

4 = 2²

OpenStudy (anonymous):

√ 12 = √ (4 * 3 ) = √ 2² * √ 3 = ...

OpenStudy (phi):

I would distribute the sqrt(6) I would factor all the numbers under the radical sign \[ \sqrt{6}(7\sqrt{12}- 10\sqrt{3})\] \[ (7\sqrt{12}\sqrt{6}- 10\sqrt{3}\sqrt{6})\] \[ (7\sqrt{12 \cdot 6}- 10\sqrt{3\cdot 6})\] now factor 12 into \(2\cdot 2 \cdot 3 \) and 6 into \(2\cdot 3 \) \[ (7\sqrt{2\cdot 2 \cdot 3 \cdot 2 \cdot 3}- 10\sqrt{3\cdot 2 \cdot 3})\] now pull out "pairs" from out of the radical \[ (7\cdot 2 \cdot 3\sqrt{2 }- 10\cdot 3\sqrt{ 2 })\] now simplify. what do you get?

OpenStudy (anonymous):

\[42\sqrt{2}-10\sqrt{3}\]